with a machine of 65% efficiency if one raises a load of 600n through a distance of 15m applying a force of effort which is equal to 600n calculate the ima(vr) of machine and the distance that an effort moved


1
Expert's answer
2021-04-26T18:00:08-0400

"\eta=\frac{M_1}{M_2}=\frac{Pl}{Fs}\implies" "l=\frac{\eta Fs}{P},"

"l=\frac{0.65\cdot 600\cdot 15}{600}=9.75~m."